eel Printing uses two measures of activity, press runs and book set-ups, in the cost formulas in its budgets and performance reports. The cost formula for wages and salaries is $8,850 per month plus $400 per press run plus $950 per book set-up. The company expected its activity in July to be 204 press runs and 111 book set-ups, but the actual activity was 201 press runs and 110 book set-ups. The actual cost for wages and salaries in July was $193,780.

The activity variance for wages and salaries in July would be closest to:

Multiple Choice

  • $2,120 F

  • $2,150 U

  • $2,120 U

  • $2,150 F
Answer #1

Answer : $ 2,150 F

Cost Formula for Wages and Salaries = 8,850+ number of Press *400+ Number of book Sheets *950

Cost at Actual Activity (Flexible Budgeted ) = 8,850+(201*400)+(110*950) = 193,750

Cost at Estimated Activity = 8,850+(204*400)+(111*950) = 195,900

Activity Varaince = Flexible budgeted Cost - Estimated activity Cost = 193750-195900 = 2,150 Favorbale